what is ln(X^(1/4)) = Sqrt(ln X)?
Answer by khwang(438)

ln(X^(1/4)) = Sqrt(ln X)
or 1/4 ln(X) = Sqrt(ln X)
or ln(X) = 4 Sqrt(ln X)
Let Ln(X) = u,
we have u = 4 sqrt(u)
Square both sides:
u^2 = 16 u,
so u(u-16) = 0
u = 0 or 16
That is, ln(X) = 0 ,X = 1 or
ln(X) = 16, X = e^16.

Please check by yourself.

Kenny